A particle of mass m moving with velocity v strike a stationary particle of mass 2m and strikes to it, the speed of the system will be 

Grade:9

1 Answers

Khimraj
5 years ago
As the external force is zero so we can apply conservation of momentum law
So
m*v = (m+2m)*x
Then x = v/3 m/s
Hope it clears. If you like answer then please approve it.